# Flat Spectrum Radio Quasars as high-energy neutrino sources

### v1

@@ 7 Discussion
Several studies aimed at the search for positional correlations between IceCube events and extragalactic sources were carried out in the last years. They all adopted different approaches and started from different neutrino events and AGN catalogues. In the following sections we summarize the main results coming from the studies.
@@ 7.1 Neutrino events
For what concerns the neutrino events, many authors (including this paper) considered the list of track-like events that offer a relatively good positional accuracy (Giommi et al. 2020a; Plavin et al. 2020, 2023), although early searches have also used cascade-type events (Padovani et al. 2016), that are characterized by larger positional uncertainties but, at the same time, a more likely astrophysical origin. Apart from the list of single neutrino events, the IceCube collaboration has also provided an all-sky map of possible neutrino excesses (Aartsen et al. 2017) in the form of a p-value, for each sky position, that gives the level of clustering in the neutrino data, i.e., a measure of the significance of neutrino events being uniformly distributed. The map is based on more than 700,000 events taken during 7 years of data collection. This map was used by the IceCube collaboration (Aartsen et al. 2017) and by other authors (Plavin et al. 2021; Buson et al. 2023; Bellenghi et al. 2023) to search for astrophysical counterparts. More recent versions of the map, based on 10 years of data, were produced and used by Aartsen et al. (2020) and Bellenghi et al. (2023).
7.2 FSRQ versus BLLac
@@
As far as the blazar catalogues are concerned, different possibilities were considered. A common choice was the selection of gamma-ray blazars, detected in one of the available FERMI catalogues (Aartsen et al. 2017, 2020) or blazars (typically BL Lacs) that are promising candidates to be detected at Very High Energy (VHE) gamma-rays (Giommi et al. 2020b). This choice was motivated by the fact that, in the hadronic model, a neutrino emission should be accompanied by the emission of a gamma-ray photon. However, not only the gamma-ray photon can be absorbed, right after its emission or during its travel towards Earth (by the diffuse extragalactic background light), but a gamma-ray emission can be also produced by other mechanisms (e.g. the Inverse Compton scattering) without the production of a neutrino (leptonic model). It is possible that both leptonic and hadronic mechanisms are at work in the jets of blazars and, therefore, a selection based on gamma-ray strength may not be the optimal choice. A gamma-ray “agnostic” approach was adopted by Buson et al. (2022, 2023) and Plavin et al. (2020, 2021, 2023) who considered, respectively, all the blazars included in the Roma-BZCat catalog (5BZCat, Massaro et al. 2015) or the radio brightest blazars, based on the Very Long Baseline Array (VLBA) calibrator surveys (VCS). In all these works, blazars are analysed as a unique class, without distinguishing between BL Lac objects and FSRQ.
@@
Overall, the most significant (3 post-trial) correlations found so far are those with the BZCat blazars (Buson et al. 2022) and with radio bright (150 mJy, Plavin et al. 2020, 2021) blazars. In all these works, the FSRQ that seem to correlate with the neutrino events are radio bright, above 200 mJy. In the “dissection” analysis discussed in Giommi et al. (2020a), a 3 (post-trials) correlation is found only with blazars (including both BL Lac and FSRQ) with the synchrotron component peaking at high frequencies (1014 Hz) while no significant correlation is found with low-frequency peaked blazars. FERMI-bright blazars do not seem to provide significant correlations except when using the most recent (10 years) map (Aartsen et al. 2020). The correlation with radio bright blazars from the RFC catalogue was confirmed at 3.3 also in the Southern Hemisphere by Bellenghi et al. (2023) using the 7-years map although nothing significant was found using the revised map based on 10-years data. This could be in part due to the lower sensitivity achieved by their analysis compared to Aartsen et al. (2020) caused by the lack of an accurate detector response matrices publicly available (Bellenghi et al. 2023).
@@
In the work presented here, we obtain a significant (4) correlation, using a sample of radio bright (100 mJy) FSRQ with declination above 20∘. The type of selection is similar to that discussed in Plavin et al. (2020) although we focus only on the FSRQ, excluding BL Lac objects. Moreover, we restrict the analysis to the sky area covered by SDSS in order to have an almost complete and uniform spectral identification of the radio bright sources. It is interesting to note that the 4 sources in the Plavin et al. (2020) that are indicated as most likely counterparts of the neutrino events are all classified as FSRQ, even if the starting sample includes both FSRQ and BL Lac objects. These four FSRQ are not included in our sample since they fall in an area of sky that we are not covering. Our result is corroborated by the observation that at high declination (Dec 20∘) the significance of the correlation is higher. Based on these results, it seems that the radio brightest FSRQ, mostly of low-frequency peak type, give a high signal in the correlation analyses carried so far.
@@
We note that the “dissection” analysis performed by Giommi et al. (2020a) did not find any significant correlation when considering low-frequency peaked blazars. This could be due to the fact that the objects considered in that work include both BL Lac and FSRQ and, more importantly, radio faint sources. For a direct comparison with our findings, this kind of analysis should be carried out considering only bright (100 mJy) FSRQ.
@@
We finally note that TXS0506+056 is formally classified as BL Lac object, and therefore a priori excluded in the analysis presented here. The fact that broad emission lines, although present (Padovani et al. 2019), were not visible in the discovery spectra is related to the fact that the synchrotron component peaks at a relatively high frequency peak (1014-1015 Hz, Padovani et al. 2018), i.e. in the visible range, which is higher compared to “normal” FSRQs, where this component typically peaks in the far infra-red band. Because of this peculiar characteristic, the non-thermal emission from the jet is able to swamp the accretion disk/BLR components, thus hiding the emission lines. These objects, also called “masquerating BL Lacs” (Giommi et al. 2013), represent a rare tail of the entire FSRQ population. Distinguishing these sources from “true” BL Lacs is not simple, but their inclusion in a future analysis might increase the significance of our results.
@@ 7.3 Jet versus accretion
As mentioned in the Introduction, neutrino production mechanisms linked to matter accretion onto the SMBH have been proposed in the literature for both NGC 1068 and TXS 0506+056. However, the results presented in this paper suggest that the neutrino production mechanism in these FSRQs is not associated with accretion. The correlation between FSRQ positions and neutrino events, in contrast with the lack of correlation of samples of QSOs with similar optical properties but no radio emission, strongly indicates that neutrino emission originates in relativistic jets. To our knowledge, it is the first time that such a direct comparison between RL and RQ QSO, in the search for counterparts of neutrino events, is carried out.
@@ 8 Summary and conclusions
Starting from an optically selected sample of AGN (Shen et al., 2011) and a well-defined, flux density-limited radio survey at 5 GHz in the Northern Hemisphere (CLASS), we identify a sample of 1122 flat-spectrum radio quasars (FSRQs), referred to as CLASH, from the larger RQ QSO population with identical optical properties. We then perform a simple positional cross-correlation between both samples and a set of 15 IceCUBE events extracted from IceCat-1, selecting events with relatively well-constrained positions within the SDSS/CLASS survey area. Using Monte Carlo randomization of positions, we estimate the p-values associated with the observed correlations. Our analysis of the CLASH sample at six different radio flux density limits revealed that the most significant correlation (p-value) occurs at a flux density limit of 100 mJy, corresponding to a post-trials significance of 2.7. In contrast, applying the same test to the RQ QSO population yielded significantly higher p-values, fully consistent with a random distribution.
Analyzing the distribution of minimum distances between neutrino events and CLASH sources, we observe a clear trend: at declinations above 20∘, these distances are systematically smaller than expected by chance, whereas closer to the celestial Equator, they tend to be larger. A Kolmogorov-Smirnov test confirms this trend. While the minimum distance distribution for the full sample differs from a random distribution at a moderate significance level of 2.2, restricting the analysis to high-declination events increases the statistical significance to 4.0. In contrast, we find that the minimum distance distribution for the RQ QSO sample is indistinguishable from a random sample. This reinforces the observed correlation between neutrinos and FSRQs.
These findings indicate that, at least some (60%) of the high energy astrophysical neutrinos detected by IceCube, may originate from blazars, particularly FSRQs, with their production likely occurring in relativistic jets rather than in the ”radio-quiet” regions, such as the accretion disk or corona. Our expectation is that most neutrino events that will be detected at high declination by IceCube, particularly those with relatively good positional accuracy, will be located close to a luminous source with a flat radio spectrum and broad emission lines in the optical.
@@ Acknowledgements.
The authors would like to thank Paolo Giommi and Lorenzo Caccianiga for useful discussions and suggestions. We acknowledge the financial support from INAF under the projects “Quasar jets in the early Universe” (Ricerca Fondamentale 2022), “Testing the obscuration in the early Universe” (Ricerca Fondamentale 2023).References
@@
- Abbasi et al. (2022) Abbasi, R., Ackermann, M., Adams, J., et al. 2022, Phys. Rev. D, 106, 022005

### v2

@@ 7 Discussion
Several studies aimed at the search for positional correlations between IceCube events and extragalactic sources were carried out in the last few years. They all adopted different approaches and started from different neutrino events and AGN catalogues. In the following sections we summarize the main results coming from the studies.
@@ 7.1 Neutrino events
Concerning neutrino events, many authors (including us) considered the list of track-like events with relatively good positional accuracies (Giommi et al. 2020a; Plavin et al. 2020, 2023).Early searches used cascade-type events (Padovani et al. 2016), which are characterized by larger positional uncertainties but have a more likely astrophysical origin. Apart from the list of single neutrino events, the IceCube collaboration has also provided an all-sky map of possible neutrino excesses (Aartsen et al. 2017) in the form of a p-value, for each sky position, that gives the level of clustering in the neutrino data, i.e. a measure of the significance of neutrino events being uniformly distributed. The map is based on more than 700,000 events taken during 7 years of data collection. This map was used by the IceCube collaboration (Aartsen et al. 2017) and by other authors (Plavin et al. 2021; Buson et al. 2023; Bellenghi et al. 2023) to search for astrophysical counterparts. More recent versions of the map, based on 10 years of data, were produced and used by Aartsen et al. (2020) and Bellenghi et al. (2023).
7.2 FSRQs versus BL Lacs
@@
As far as the blazar catalogues are concerned, different possibilities were considered. A common choice was the selection of gamma-ray blazars, detected in one of the available Fermi catalogues (Aartsen et al. 2017, 2020) or blazars (typically BL Lacs) that are promising candidates to be detected at very high-energy gamma rays (Giommi et al. 2020b). This choice was motivated by the fact that, in the hadronic model, a neutrino emission should be accompanied by the emission of a gamma-ray photon. However, not only the gamma-ray photon can be absorbed, right after its emission or during its travel towards Earth (by the diffuse extragalactic background light), but a gamma-ray emission can be also produced by other mechanisms (e.g. the inverse Compton scattering) without the production of a neutrino (leptonic model). It is possible that both leptonic and hadronic mechanisms are at work in the jets of blazars, and therefore, a selection based on gamma-ray strength may not be the optimal choice. A gamma-ray ‘agnostic’ approach was adopted by Buson et al. (2022, 2023) and Plavin et al. (2020, 2021, 2023) who considered, respectively, all the blazars included in the Roma-BZCat catalogue (5BZCat; Massaro et al. 2015) and the radio brightest blazars, based on the Very Long Baseline Array calibrator surveys. In all these works, blazars are analysed as a unique class, without distinguishing between BL Lac objects and FSRQs.
@@
Overall, the most significant (3 post-trial) correlations found so far are those with the BZCat blazars (Buson et al. 2022) and with radio bright (150 mJy, Plavin et al. 2020, 2021) blazars. In all these works, the FSRQs that seem to correlate with the neutrino events are radio bright, above 200 mJy. In the ‘dissection’ analysis discussed in Giommi et al. (2020a), a 3 (post-trials) correlation is found only with blazars (including both BL Lacs and FSRQs) with the synchrotron component peaking at high frequencies (1014 Hz) while no significant correlation is found with low-frequency peaked blazars. Fermi-bright blazars do not seem to provide significant correlations except when using the most recent (10 years) map (Aartsen et al. 2020). The correlation with radio bright blazars from the RFC catalogue (Petrov & Kovalev, 2025) was confirmed at 3.3 also in the southern hemisphere by Bellenghi et al. (2023) using the 7-year map although nothing significant was found using the revised map based on 10 years of data. This could be in part due to the lower sensitivity achieved by their analysis compared to Aartsen et al. (2020) caused by the lack of an accurate detector response matrices publicly available (Bellenghi et al. 2023).
@@
In the work presented here, we obtain a significant correlation (4) based on a sample of radio bright (100 mJy) FSRQs with declinations above 20∘. The type of selection is similar to that discussed in Plavin et al. (2020) although we focused only on the FSRQs, excluding BL Lac objects. Moreover, we restricted the analysis to the sky area covered by SDSS in order to have an almost complete and uniform spectral identification of the radio bright sources. It is interesting to note that the four sources in the Plavin et al. (2020) that are indicated as most likely counterparts of the neutrino events are all classified as FSRQs, even if the starting sample includes both FSRQs and BL Lac objects. These four FSRQs are not included in our sample since they fall in an area of sky that we are not covering. Our result is corroborated by the observation that at high declination (Dec. 20∘) the significance of the correlation is higher. Based on these results, it seems that the radio-brightest FSRQs, mostly of low-frequency peak type, give a high signal in the correlation analyses carried so far.
@@
We note that the ‘dissection’ analysis performed by Giommi et al. (2020a) did not find any significant correlation when considering low-frequency peaked blazars. This could be due to the fact that the objects considered in that work include both BL Lacs and FSRQs and, more importantly, radio faint sources. For a direct comparison with our findings, this kind of analysis should only be carried out on bright FSRQs (100 mJy).
@@
We finally note that TXS0506+056 is formally classified as a BL Lac object and, therefore, a priori excluded in the analysis presented here. The fact that broad emission lines, although present (Padovani et al. 2019), were not visible in the discovery spectra is related to the fact that the synchrotron component peaks at a relatively high frequency peak (1014-1015 Hz, Padovani et al. 2018), i.e. in the visible range, which is higher compared to ‘normal’ FSRQs, where this component typically peaks in the far infra-red band. Because of this peculiar characteristic, the non-thermal emission from the jet is able to swamp the accretion disk and BLR components, thus hiding the emission lines. These objects, also called ‘masquerading BL Lacs’ (Giommi et al. 2013), represent a rare tail of the entire FSRQ population. Distinguishing these sources from ‘true’ BL Lacs is not simple, but their inclusion in a future analysis might increase the significance of our results.
@@ 7.3 Jet versus accretion
As mentioned in the Introduction, neutrino production mechanisms linked to matter accretion onto the SMBH have been proposed in the literature for both NGC 1068 and TXS 0506+056. However, the results presented in this paper suggest that the neutrino production mechanism in these FSRQs is not associated with accretion. The correlation between FSRQ positions and neutrino events, in contrast with the lack of correlation of samples of quasars with similar optical properties but no radio emission, strongly indicates that neutrino emission originates in relativistic jets. To our knowledge, it is the first time that such a direct comparison between radio-loud and RQ quasars, in the search for counterparts of neutrino events, is carried out.
@@ 8 Summary and conclusions
Starting from an optically selected sample of AGNs (Shen et al., 2011) and a well-defined, flux-density-limited radio survey at 5 GHz in the northern hemisphere (CLASS), we identified a sample of 1122 FSRQs, referred to as the CLASH sample, from the larger RQ quasar population with identical optical properties. We then performed a simple positional cross-correlation between both samples and a set of 15 IceCube events extracted from IceCat-1, selecting events with relatively well-constrained positions within the SDSS/CLASS survey area. Using a Monte Carlo randomization of positions, we estimated the p-values of the observed correlations. Our analysis of the CLASH sample at six different radio flux density limits reveals that the most significant correlation occurs at a flux density limit of 100 mJy, corresponding to a post-trial significance of 2.7. In contrast, applying the same test to the RQ quasar population yielded significantly higher p-values that were fully consistent with a random distribution.
Analysing the distribution of minimum distances between neutrino events and CLASH sources, we observe a clear trend: at declinations above 20∘, these distances are systematically smaller than what would be expected by chance, whereas closer to the celestial equator, they tend to be larger. A KS test confirms this trend. While the minimum distance distribution for the full sample differs from a random distribution at a moderate significance level of 2.2, restricting the analysis to high-declination events increases the statistical significance to 4.0. In contrast, we find that the minimum distance distribution for the RQ quasar sample is indistinguishable from a random sample. This reinforces the observed correlation between neutrinos and FSRQs.
These findings indicate that a majority (60%) of the high-energy astrophysical neutrinos detected by IceCube originate from blazars, particularly FSRQs. Their production likely occurs in relativistic jets rather than in the RQ regions, such as the accretion disk or corona. Our expectation is that most neutrino events that will be detected at high declinations by IceCube, particularly those with relatively good positional accuracies, will be located close to a luminous source with a flat radio spectrum and broad emission lines in the optical.
@@ Acknowledgements.
The authors thank Paolo Giommi and Lorenzo Caccianiga for valuable discussions and suggestions. We acknowledge the financial support from INAF under the projects “Quasar jets in the early Universe” (Ricerca Fondamentale 2022) and “Testing the obscuration in the early Universe” (Ricerca Fondamentale 2023).References
@@
- Abbasi et al. (2022a) Abbasi, R., Ackermann, M., Adams, J., et al. 2022a, Phys. Rev. D, 106, 022005
- Abbasi et al. (2022b) Abbasi, R., Ackermann, M., Adams, J., et al. 2022b, ApJ, 938, 38
@@
- Murase (2022) Murase, K. 2022, ApJ, 941, L17
- Murase et al. (2014) Murase, K., Inoue, Y., & Dermer, C. D. 2014, Phys. Rev. D, 90, 023007
@@
- Petrov & Kovalev (2025) Petrov, L. & Kovalev, Y. 2025, in American Astronomical Society Meeting Abstracts, Vol. 245, American Astronomical Society Meeting Abstracts #245, 448.03